OpenStudy (anonymous):

Find the exact length of the curve y=ln((1+e^(-x))/(1-e^(-x))) between x=1 and x=2

OpenStudy (dumbcow):

\[Length = \int\limits_{a}^{b}\sqrt{1+f'(x)^{2}} dx\]

OpenStudy (anonymous):

the derivative on wolfram is 2e^x/e^(2x)-1

OpenStudy (dumbcow):

i get f'(x) \[f'(x) = \frac{-2e^{-x}}{(1+e^{-x})(1-e^{-x})}\]

OpenStudy (dumbcow):

oh ok, well i didn't simplify it enough

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\sqrt{1+4e^(2x)/(e^(2x)-1)^2}\]

OpenStudy (anonymous):

i dont know how to integrate that

OpenStudy (zarkon):

I haven't worked it out...but you should probably combine to make one term under the radical

OpenStudy (zarkon):

what I wrote above is what you want to do